# MedualTime: A Dual-Adapter Language Model for Medical Time Series-Text Multimodal Learning

### v3

@@ Abstract
The recent rapid advancements in language models (LMs) have garnered attention in medical time series-text multimodal learning. However, existing contrastive learning-based and prompt-based LM approaches tend to be biased, often assigning a primary role to time series modality while treating text modality as secondary. We classify these approaches under a temporal-primary paradigm, which may overlook the unique and critical task-relevant information embedded in text modality like clinical reports, thus failing to fully leverage mutual benefits and complementarity of different modalities. To fill this gap, we propose a novel textual-temporal multimodal learning paradigm that enables either modality to serve as the primary while being enhanced by the other, thereby effectively capturing modality-specific information and fostering cross-modal interaction. In specific, we design MedualTime, a language model composed of dual adapters to implement temporal-primary and textual-primary modeling simultaneously. Within each adapter, lightweight adaptation tokens are injected into the top layers of LM to encourage high-level modality fusion. The shared LM pipeline by dual adapters not only achieves adapter alignment but also enables efficient fine-tuning, reducing computational resources. Empirically, MedualTime demonstrates superior performance on medical data, achieving notable improvements of 8% accuracy and 12% F1 in supervised settings. Furthermore, MedualTime’s transferability is validated by few-shot label transfer experiments from coarse-grained to fine-grained medical data. https://github.com/start2020/MedualTime
footnotetext: Co-first authors, Corresponding authors.
@@ 1 Introduction
  • MedualTime demonstrates superior performance on public medical multimodal datasets, showing its strong generalization and transferability in medical diagnosis tasks. Notably, it achieves an average improvement of 8% in accuracy and 12% in F1 score under supervised learning.
@@ 2 Related Work
All these LM-based multimodal works consider time series as the primary modality for decision-making, with text serving as an auxiliary modality to enhance time series modeling. In contrast, MedualTime allows each modality to act as a primary modality, which can comprehensively capture the unique and shared semantics provided by different modalities.
@@ 3.1 Textual-primary Multimodal Adapter
Processed by the textual tokenizer, the text input can be modeled by $I^{s}$-length word tokens with embedding $\bm{E}_{s}\in\mathbb{R}^{I^{s}\times D}$, where $D$ is the hidden dimension. For the first $L-M$ transformer layers, which are standard transformer layers, the forward process of layer-$l$ is:
@@ 3.4 Training Loss
[…]
@@ 4 Classes 5 Classes 2 Classes 4 Classes Average
Time + Text MedualTime 0.83 0.78 0.75 0.76 0.81 0.75 0.75 0.75 0.87 0.75 0.65 0.68 0.79 0.82 0.74 0.75 0.83 0.73
@@ 4.1 Experimental Setup
Datasets Since our approach focuses on medical tasks involving coupled time-series and text modalities, where each modality provides valuable information for decision-making, we conduct experiments using publicly available datasets that meet the necessary criteria above: (1) PTB-XL https://physionet.org/content/ptb-xl/1.0.3/ Wagner et al. (2020): This electrocardiogram (ECG) corpus consists of 12-lead ECG signals, which capture the electrical activity of the heart, along with clinical reports describing signal characteristics without diagnostic labels. PTB-XL provides two label sets: a coarse-grained label set for disease detection (four classes) and a fine-grained label set for specific disease classification (five classes). (2) TUSZ v1.5.2 https://isip.piconepress.com/projects/nedc/html/tuh_eeg/ Shah et al. (2018): The Temple University Seizure Corpus (TUSZ) is a large-scale corpus of EEG (electroencephalogram) signals that record the electrical activity of the brain. It includes 19-channel EEG recordings and the clinical history for each patient session. Similar to PTB-XL, TUSZ offers two label sets: a coarse-grained label set for distinguishing seizure and non-seizure EEG signals (two classes), and a fine-grained label set for seizure type classification (four classes). More details about the datasets, including the label sets, data splits, and preprocessing steps, are provided in Appendix 1.1.
@@
Implementations Inspired by Liu et al. (2024); Jia et al. (2024); Zhou et al. (2024), which have validated GPT-2’s capability in handling time-series data, MedualTime adopts a frozen GPT-2 as the backbone. In the textual-primary multimodal adapter, the tokenizer is derived from GPT-2. To avoid heavy computational costs, we use a lightweight CNN-based model as the temporal encoder, consisting of three convolutional blocks, each with three CNN layers. This encoder is trained from scratch to adapt to the specific tasks. In the temporal-primary multimodal adapter, following METS Li et al. (2024), a frozen ClinicalBERT Wang et al. (2023) pre-trained on medical data as the textual encoder. Ablation study about the backbone and textual encoder is in Appendix 1.5. All hidden dimensions are set to 768 to match the backbone (i.e., GPT-2). The time series patch size and stride are both set to 25. Adam is adopted as the optimizer Kingma (2014). All experiments are implemented by PyTorch Framework with NVIDIA A6000 (48G) GPU.
@@ 4.2 Supervised Learning
Figure 4: Few-shot experiment results on accuracy. The red dashed line is MedualTime.
@@ 4 Classes 5 Classes 2 Classes 4 Classes Average
Time + Text MedualTime 0.77 0.65 0.77 0.58 0.76 0.57 0.74 0.62 0.76 0.61
@@ 4.4 Unsupervised Learning
We follow the experimental setting in previous works Zhang et al. (2023b) to assess our model’s ability to generate general representations without ground truth supervision. First, we conduct unsupervised experiments. Then, using the unsupervised embeddings obtained, we train a linear classifier with varying proportions (from 10% to 100%) of labeled training data. Table 2 shows the results for the 100% labeled training data proportion. More detailed results, analysis, and visualization can be found in Appendix 1.4.
@@ 4 Classes 5 Classes 2 Classes 4 Classes
Figure 5: Embedding visualizations of different encoders on PTB-XL (4 Classes), with labels distinguished by color.
@@
(e) Additional Experiments Due to space limitations, additional experiments on model design are presented in the Appendix 1.5, including: (1) Backbone Analysis to further demonstrate MedualTime’s generalization across different language models, (2) Modality Fusion Mechanism Analysis to compare our adapter fusion approach with other methods such as simple concatenation, weighted fusion, and attention mechanisms. (3) Sensitive Analysis to evaluate the impact of two key hyperparameters (multimodal fusion layers M and the adaptation token length P) in Appendix 1.6.

### v4

@@ Abstract
The recent rapid advancements in language models (LMs) have garnered attention in medical time series-text multimodal learning. However, existing contrastive learning-based and prompt-based LM approaches tend to be biased, often assigning a primary role to time series modality while treating text modality as secondary. We classify these approaches under a temporal-primary paradigm, which may overlook the unique and critical task-relevant information embedded in text modality like clinical reports, thus failing to fully leverage mutual benefits and complementarity of different modalities. To fill this gap, we propose a novel textual-temporal multimodal learning paradigm that enables either modality to serve as the primary while being enhanced by the other, thereby effectively capturing modality-specific information and fostering cross-modal interaction. In specific, we design MedualTime, a language model composed of dual adapters to implement temporal-primary and textual-primary modeling simultaneously. Within each adapter, lightweight adaptation tokens are injected into the top layers of LM to encourage high-level modality fusion. The shared LM pipeline by dual adapters not only achieves adapter alignment but also enables efficient fine-tuning, reducing computational resources. Empirically, MedualTime demonstrates superior performance on medical data, achieving notable improvements of 8% accuracy and 12% F1 in supervised settings. Furthermore, MedualTime’s transferability is validated by few-shot transfer experiments from coarse-grained to fine-grained medical data.
@@ 1 Introduction
  • MedualTime demonstrates superior performance on public medical multimodal datasets, showing its strong generalization and transferability in medical diagnosis tasks. Notably, it achieves an average improvement of 8% in accuracy and 12% in F1 score under supervised learning. The code is in https://github.com/start2020/MedualTime.
@@ 2 Related Work
All these LM-based multimodal works consider time series as the primary modality for decision-making, with text serving as an auxiliary modality to enhance time series modeling. In contrast, MedualTime allows each modality to act as a primary modality, which can comprehensively capture the unique and shared semantics in different modalities.
@@ 3.1 Textual-primary Multimodal Adapter
Processed by the textual tokenizer, the text input can be modeled by $I^{s}$-length word tokens with embedding $\bm{E}_{s}\in\mathbb{R}^{I^{s}\times D}$, where $D$ is the hidden dimension. For the first LML-M transformer layers, which are standard transformer layers, the forward process of layer-$l$ is:
@@ 3.4 Training Loss
$\begin{aligned} \mathcal{L}_{cross}=-\sum_{i=1}^{B}\bigg{(}&\log\frac{\exp\left(\operatorname{sim}\left(\bm{H}_{s,i}^{L},\bm{H}_{t,i}^{L}\right)/\tau\right)}{\sum_{k=1}^{B}\mathds{1}_{[k\neq i]}\exp\left(\operatorname{sim}\left(\bm{H}_{s,i}^{L},\bm{H}_{t,k}^{L}\right)/\tau\right)}\\ &+\log\frac{\exp\left(\operatorname{sim}\left(\bm{H}_{t,i}^{L},\bm{H}_{s,i}^{L}\right)/\tau\right)}{\sum_{k=1}^{B}\mathds{1}_{[k\neq i]}\exp\left(\operatorname{sim}\left(\bm{H}_{t,i}^{L},\bm{H}_{s,k}^{L}\right)/\tau\right)}\bigg{)}\end{aligned}$
@@ 4 Classes 5 Classes 2 Classes 4 Classes Average
Time
Text
MedualTime 0.83 0.78 0.75 0.76 0.81 0.75 0.75 0.75 0.87 0.75 0.65 0.68 0.79 0.82 0.74 0.75 0.83 0.73
@@ 4.1 Experimental Setup
Datasets Since our approach focuses on medical tasks involving coupled time-series and text modalities, where each modality provides valuable information for decision-making, we conduct experiments using publicly available datasets that meet the necessary criteria above: (1) PTB-XL  [^1] https://physionet.org/content/ptb-xl/1.0.3/ Wagner et al. (2020): This electrocardiogram (ECG) corpus consists of 12-lead ECG signals, which capture the electrical activity of the heart, along with clinical reports describing signal characteristics without diagnostic labels. PTB-XL provides two label sets: a coarse-grained label set for disease detection (four classes) and a fine-grained label set for specific disease classification (five classes). (2) TUSZ v1.5.2  [^2] https://isip.piconepress.com/projects/nedc/html/tuh_eeg/ Shah et al. (2018): The Temple University Seizure Corpus (TUSZ) is a large-scale corpus of EEG (electroencephalogram) signals that record the electrical activity of the brain. It includes 19-channel EEG recordings and the clinical history for each patient session. Similar to PTB-XL, TUSZ offers two label sets: a coarse-grained label set for distinguishing seizure and non-seizure EEG signals (two classes), and a fine-grained label set for seizure type classification (four classes). More details about the datasets, including the label sets, data splits, and preprocessing steps, are provided in Appendix 1.1.
@@
Implementations Inspired by Liu et al. (2024); Jia et al. (2024); Zhou et al. (2024), which have validated GPT-2’s capability in handling time-series data, MedualTime adopts a frozen GPT-2 as the backbone. In the textual-primary multimodal adapter, the tokenizer is derived from GPT-2. To avoid heavy computational costs, we use a lightweight CNN-based model as the temporal encoder, consisting of three convolutional blocks, each with three CNN layers. This encoder is trained from scratch to adapt to the specific tasks. In the temporal-primary multimodal adapter, following METS Li et al. (2024), a frozen ClinicalBERT Wang et al. (2023) pre-trained on medical data as the textual encoder. All hidden dimensions are set to 768 to match the backbone (i.e., GPT-2). The time series patch size and stride are both set to 25. Adam is adopted as the optimizer Kingma (2014). All experiments are implemented by PyTorch Framework with NVIDIA A6000 (48G) GPU.
@@ 4.2 Supervised Learning
Figure 4: Few-shot experiment results on accuracy. The green line is MedualTime.
@@ 4 Classes 5 Classes 2 Classes 4 Classes Average
Time
Text
MedualTime 0.77 0.65 0.77 0.58 0.76 0.57 0.74 0.62 0.76 0.61
@@ 4.4 Unsupervised Learning
We follow the experimental setting in previous works Zhang et al. (2023b) to assess our model’s ability to generate general representations without ground truth supervision. First, we conduct unsupervised experiments. Then, using the unsupervised embeddings obtained, we train a linear classifier with varying proportions (from 10% to 100%) of labeled training data. Table 2 shows the results for the 100% labeled training data proportion.
@@ 4 Classes 5 Classes 2 Classes 4 Classes
Figure 5: Embedding visualizations of different encoders on PTB-XL (4 Classes), i.e. ClinicalBERT, TS2Vec, MedualTime (Text), MedualTime (Time), MedualTime, from left to right. Different colors represent different labels.
@@ Acknowledgments
Contribution Statement
Co-first Authors, Corresponding Authors.